OpenStudy (anonymous):

HELP??? reciprocal of -2/15? I got -1/30, is this correct?

OpenStudy (anonymous):

reciprocal of -2/15 = \(\dfrac{1}{\frac{-2}{15}}\)

OpenStudy (anonymous):

\[\dfrac{1}{\frac{-2}{15}}\\~\\~\\\dfrac{1}{\frac{-2}{15}}\cdot1\\~\\~\\\dfrac{1}{\frac{-2}{15}}\cdot\dfrac{15}{15}\\~\\~\\\dfrac{15}{-2}\]

OpenStudy (anonymous):

makes sense, thank you!

OpenStudy (anonymous):

is 15 negative as well?

OpenStudy (anonymous):

or just the 2

OpenStudy (anonymous):

-(2/15)

OpenStudy (anonymous):

well... 15 could be negative, but then 2 must be positive. \(\dfrac{15}{-2}=-\dfrac{15}{2}=\dfrac{-15}{2}\)

OpenStudy (anonymous):

ok, i understand. can you maybe help me with one more question?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

-(9/8)/(3/5)

OpenStudy (anonymous):

I got 1 7/8 on my calculator.

OpenStudy (anonymous):

Almost, it's \(-1\dfrac{7}{8}\)

OpenStudy (anonymous):

you just forgot negative sign

OpenStudy (anonymous):

oh yeah, thats what was on my calculator. thanks!

OpenStudy (anonymous):

i'd give you another medal if i could.

OpenStudy (anonymous):

lol it's ok. you are welcome :)

OpenStudy (texaschic101):

with the reciprocal...just flip the number, but keep the sign. -2/3...reciprocal = -3/2 78....reciprocal = 8/7 and if you have a whole number, such as 3, it is the same as 3/1, so the reciprocal would be 1/3. and when your dividing fractions....just flip what you are dividing by, then multiply (-9/8) / (3/5).....flip what you are dividing by, then multiply (-9/8) * (5/3) = -45/24...and when turned to a mixed number, you get - 1 7/8 make sure you flip the right number.....becuse if you do this... (-9/8) / (3/5) = (-8/9) * (3/5).....this would be wrong...you have to flip what you are dividing by....not the other number.

OpenStudy (anonymous):

Thanks @texaschic101 and sorry for the late reply but you were really helpful. I would give you a medal as well, but I already gave micahwood50 the best response, so I can't. /: Maybe next time? Thanks again! :)